not sure if this is the right section, but here it goes:
A complete algebraic solution to explain your reasoning is required to achieve full marks. Show all your work.
Write an equation in the form y = ax^2 + bx + c for the quadratic function whose graph passes through the points (8,0), (0,8), and (-2,0).

This is Grade 11 University Math, it is a TIPS problem that we got to take home but I'll have to re-write it in class with a different but similar question.

Answer by stanbon(75887)

Substitute your point values into the equation to get:
64a+8b+c=0
0a+0b+c=8
4a-2b+c=0
Solve this system of equations for a,b,and c
using, elimination, matrices, or Cramer's method.
Results are a=-.5 b=3 c=8
Cheers,
Stan H.